OpenStudy (anonymous):

Solve the system of equations by the substitution method. If the system does not have a single ordered pair as a solution, state whether the system is inconsistent or dependent. x+3y=32-3x+2y=3

OpenStudy (anonymous):

i have A (4,10), B no solution; inconsistent, C (5,9), D (-5,10)

OpenStudy (anonymous):

plz plz help me

OpenStudy (anonymous):

\[\{x+3 y=32,2 y-3 x=3\} \](5,9) if the above are the equations.

OpenStudy (anonymous):

ty very much!!!

OpenStudy (anonymous):

its written out just like this x+3y=32-3x+2y=3

OpenStudy (anonymous):

is it still (5,9) ?

OpenStudy (anonymous):

The problem expression needs a comma somewhere to make sense. A comma following the number 32 seemed to make the most sense. You have to have two equations to find the two unknowns, x and y.

OpenStudy (anonymous):

Solve the system of equations graphically. If the system does not have a single ordered pair as a solution, state whether the system is inconsistent of dependent. 9x - 2y = 30 5x + 4/3y = 24

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

answers r A( 4,3), B (3,1), C (4,-3) ,D (3,4)

OpenStudy (anonymous):

the answer is (5,9)